the length of a rectangle is 4 less than 3 times the width. The area is 160 square inches. Find the permeter (so first, find the width and length)
Answer by checkley77(12844)

L=3W-4
LW=160
(3W-4)W=160
3W^2-4W-160=0
(3W+20)(W-8)=0
W-8=0
W=8 ANS. FOR THE WIDTH.
L=3*8-4
L=24-4
L=20 ANS. FOR THE LENGTH.
PROOF:
8*20=160
160=160